# Enhancing Transformers Through Conditioned Embedded Tokens

### v1

@@ Abstract
Figure 1: Schematic representation of conditioned embeddings for a vision transformer: An image is divided into $N$ patches, with each patch then embedded as a high-dimensional vector $X_{i}\in\mathbb{R}^{d}$ for $1\leq i\leq N$. These vectors are then concatenated to form the embedded token matrix $X=[X_{1}\cdots X_{N}]^{T}$. To improve the conditioning of the self-attention mechanism, a correction term $C=[C_{1}\cdots C_{N}]^{T}$ is added to $X$, reducing its condition number. The modified matrix is then fed into the first layer of the transformer (positional encoding not shown). Its effect on the self-attention equation is illustrated in the green equation (for simplicity, layer normalization has been omitted).
@@ 2 Related Work
Conditioning: In [17], the authors investigate the condition number of weight matrices in feedforward neural networks and demonstrate that lower condition numbers lead to improved accuracy across various applications. They propose a preconditioner matrix that multiplies each layer’s weight matrix, effectively reducing its condition number and enhancing the training of dense weights. In [12], neural network optimization is explored through the lens of the neural tangent kernel. The authors introduce the concept of a PL* condition—a variant of the Polyak-Lojasiewicz condition—and show that this approach improves the conditioning of the neural tangent kernel, which, in the infinite-width limit, governs the dynamics of gradient descent [10]. Finally, [1] demonstrates that increasing the depth of feedforward neural networks improves their conditioning, thereby leading to more efficient optimization.
@@ 3 Notation
The matrix $[Ex_{1}\cdots Ex_{n}]^{T}\in\mathbb{R}^{N\times d}$ represents the embedded tokens associated to the input data. To capture the order of the tokens, a positional encoding matrix $P\in\mathbb{R}^{N\times d}$ is added forming $X=[Ex_{1}\cdots Ex_{n}]^{T}+P\in\mathbb{R}^{N\times d}$. The matrix $X$ is then input into the first layer of the transformer. For the theoretical Sec. 4 we will often use the notation $X$ to denote embedded tokens often leaving out explicit mention of the positional encoding part for ease of notation.
@@
$[\mathbf{A}_{1},\cdots,\mathbf{A}_{h}],$ (6)
@@ 5.3 Language Models
We conducted a second experiment training a GPT-2 model with a decoder-only transformer architecture, consisting of 12 layers, 12 attention heads per layer, and an embedding dimension of 512. The model was trained using masked self-attention to predict the next token, leveraging the TinyStories dataset [6] for language generation. Two versions of the model were trained from scratch: one using the original embedded tokens as in [6] and another incorporating the conditioned embedded tokens from Sec. 4.2. Both models were evaluated using perplexity loss (PPL), a widely used metric for assessing language models on datasets like TinyStories. Perplexity measures the model’s confidence in its predictions by computing the exponential of the average negative log-likelihood of the predicted tokens. A lower perplexity score indicates that the model assigns higher probabilities to correct tokens, leading to better fluency and coherence in text generation.
Val loss PPL loss
GPT-2 original 2.41 11.2
GPT-2 conditioned (ours) 2.36 10.1
Table 4: GPT-2 model trained on the TinyStories dataset. We compare two models, the original model and one using conditioned embedded tokens (conditioned). We report the validation loss and perplexity loss (PPL). In both cases our model yields a lower loss showing better performance.
@@
Results and Analysis: Tab. 4 presents the results of our experiment, demonstrating that the model trained with conditioned embedded tokens consistently achieves lower validation loss and lower perplexity compared to the original GPT-2 model. These results indicate that conditioning the embedded tokens improves model performance, leading to more accurate language generation. Fig. 3 plots the condition numbers of the embedded tokens, self-attention in first layer averaged over all heads and self-attention averaged over all layers. Each of these are evaluated over each training epoch and then averaged over all epochs.
@@
Nyströmformer conditioned (ours) 37.9 64.9 80.9 40.1 80.0
Table 5: Comparison of a Nyströmformer using its original embedded tokens (original) with a Nyströmformer using conditioned embedded tokens (ours) on the LRA benchmark. We report the evaluation accuracy (%). As can be seen from the table our methodology yields a hihger accuracy on each task.

### v2

@@ Abstract
Figure 1: Schematic representation of conditioned embeddings for a vision transformer: An image is divided into NN patches, with each patch then embedded as a high-dimensional vector $X_{i}\in\mathbb{R}^{d}$ for $1\leq i\leq N$. These vectors are then concatenated to form the embedded token matrix $X=[X_{1}\cdots X_{N}]^{T}$. To improve the conditioning of the self-attention mechanism, a correction term $C=[C_{1}\cdots C_{N}]^{T}$ is added to $X$, reducing its condition number. The modified matrix is then fed into the first layer of the transformer (positional encoding not shown). Its effect on the self-attention equation is illustrated in the green equation (for simplicity, layer normalization has been omitted).
@@ 2 Related Work
Conditioning: In [32], the authors investigate the condition number of weight matrices in feedforward neural networks and demonstrate that lower condition numbers lead to improved accuracy across various applications. They propose a preconditioner matrix that multiplies each layer’s weight matrix, effectively reducing its condition number and enhancing the training of dense weights. In [19], neural network optimization is explored through the lens of the neural tangent kernel. The authors introduce the concept of a PL* condition, a variant of the Polyak-Lojasiewicz condition, and show that this approach improves the conditioning of the neural tangent kernel, which, in the infinite-width limit, governs the dynamics of gradient descent [15]. For feedforward networks, Agarwal et al. [1] showed that increasing depth improves conditioning and thereby facilitates optimization, while MacDonald et al. [21] demonstrated that normalization can further enhance conditioning. In the context of implicit neural representations (INRs), several works have established that activation choice strongly influences conditioning and optimization performance [27, 6, 26, 12, 28, 7, 24, 29]. For transformers, skip connections have been shown to improve the conditioning of the attention block [17], while multi-head attention and polynomial activations have been shown to condition the Jacobian of the attention mechanism more effectively [31, 30]. Finally, in fine-tuning settings, conditioning adapter modules has been found to yield faster and more stable convergence [16, 3, 2].
@@ 3 Notation
The matrix $Ex_{1}\cdots Ex_{n}]^{T}\in\mathbb{R}^{N\times d}$ represents the embedded tokens associated to the input data. To capture the order of the tokens, a positional encoding matrix $P\in\mathbb{R}^{N\times d}$ is added forming $X=[Ex_{1}\cdots Ex_{n}]^{T}+P\in\mathbb{R}^{N\times d}$. The matrix $X$ is then input into the first layer of the transformer. For the theoretical Sec. 4 we will often use the notation $X$ to denote embedded tokens often leaving out explicit mention of the positional encoding part for ease of notation.
@@
$\mathbf{A}_{1},\cdots,\mathbf{A}_{h}],$ (6)
@@ 5.3 Language Models
We conducted a second experiment training a GPT-2 model with a decoder-only transformer architecture, consisting of 12 layers, 12 attention heads per layer, and an embedding dimension of 512. The model was trained using masked self-attention to predict the next token, leveraging the TinyStories dataset [10] for language generation. Two versions of the model were trained from scratch: one using the original embedded tokens as in [10] and another incorporating the conditioned embedded tokens from Sec. 4.2. Both models were evaluated using a validation loss.
Val loss
GPT-2 original 2.41
GPT-2 conditioned (ours) 2.36
Table 4: GPT-2 model trained on the TinyStories dataset. We compare two models, the original model and one using conditioned embedded tokens (conditioned). We report the validation loss, with the conditioned model achieving lower loss and thus better performance.
@@
Results and Analysis: Tab. 4 presents the results of our experiment, demonstrating that the model trained with conditioned embedded tokens consistently achieves lower validation loss when compared to the original GPT-2 model. These results indicate that conditioning the embedded tokens improves model performance, leading to more accurate language generation. Fig. 3 plots the condition numbers of the embedded tokens, self-attention in first layer averaged over all heads and self-attention averaged over all layers. Each of these are evaluated over each training epoch and then averaged over all epochs.
@@
Nyströmformer conditioned (ours) 37.9 64.9 80.9 40.1 73.3
Table 5: Comparison of a Nyströmformer using its original embedded tokens (original) with a Nyströmformer using conditioned embedded tokens (ours) on the LRA benchmark. We report the evaluation accuracy (%). As can be seen from the table our methodology yields a higher accuracy on each task.